OpenStudy (anonymous):

Determine if the function h(x) = 2 – x^2 + 3x^4 is even, odd, or neither. Justify your answer.

OpenStudy (anonymous):

a polynominal with all even exponents is even

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

however, you are supposed to write \[h(-x)=2-(-x)^2+3(-x)^4=2-x^3+3x^4=h(x)\] therefore even since \(h(-x)=h(x)\)